# Security Concerns for Large Language Models: A Survey

### v2

@@ III-C Recent Empirical Studies and Advanced Techniques
TABLE II: Summary of Adversarial Attacks on LLMs
@@
Inference-Time Input Perturbations Minor changes to input (synonyms, misspellings, token manipulation) to cause misclassification or harmful output. Evades simple filters, exploits model sensitivity to token sequences. Zou et al. (GCG) [7] (transferable suffixes), Chen et al. (Imprompter) [21] (steganographic data extraction).
Inference-Time Universal Triggers Carefully chosen rare token sequences in prompts to induce specific outputs. Robust, transferable across prompts. Wallace et al. [26] (GPT-2 outputs attacker-specified content).
Inference-Time Fuzzing Automated discovery of subtle token-level changes that flip LLM behavior. Exposes brittleness to minor input variations (whitespace, emojis). FuzzSurvey [27] (Many chat-LLMs brittle to small input changes).
IV Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Because LLMs excel at generating fluent, coherent, and contextually appropriate text, they can be readily misused to automate and scale social engineering attacks and various forms of cybercrime. The range of such misapplications is broad, leveraging the LLMs’ generative prowess for nefarious ends. An overview of these misuse categories, the specific LLM capabilities they exploit, and prominent examples are detailed in Table III.
IV-A Automating Social Engineering and Cybercrime
@@
The Imprompter proposed by Chen et al. [21] further illustrates that steganographically encoded inputs can be used to embed malicious instructions in benign-looking content, enabling stealthy exfiltration and phishing attempts via covert LLM channels.
@@
IV-B Generation of Disinformation and Deceptive Content
@@
IV-C Emergence of Specialized Malicious LLMs and Ecosystem Exploitation
@@
TABLE III: Summary of LLM Misuse by Malicious Actors
@@
Phishing Social Engineering Generating convincing phishing emails, tailored spam, devising fraud strategies. Fluent text generation, contextual understanding, mimicry. Saha Roy et al. [8], Cohen et al. (Morris-II) [18], Chen et al. (Imprompter) [21] Scalable, automated, convincing attacks; autonomous email worms.
@@
Specialized Malicious LLMs Custom ”jailbroken” or fine-tuned models for malicious tasks. Transfer learning, fine-tuning capabilities. WormGPT, FraudGPT [29], PoisonGPT [30], Wan et al. [31] Democratization of advanced malicious tools.
@@
V Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools or interact with external environments, novel categories of risks emerge. These risks stem not just from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which may not align with those of human designers or users [34], as extensively detailed in discussions of catastrophic risks from agentic AI [35]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table IV provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
V-A Goal Misalignment and Unfaithful Reasoning
@@
Modern LLMs often employ Chain-of-Thought (CoT) prompting to produce step-by-step reasoning before an answer. Some may believe that through careful monitoring the CoT, we can gain insight into the model’s internal decision-making process and detect early signs of misalignment. However, Lanham et al. [36] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study,” which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[17] demonstrated that even if CoTs can be monitored for misbehaviors (e.g., reward hacking), directly penalizing “bad thoughts” by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking”. In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4 released in May 2025, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior,” indicating “moderate overall reasoning faithfulness” [37].
TABLE IV: Summary of Intrinsic Risks in Autonomous LLM Agents
@@
V-B Emergent Deception and Self-Preservation Behaviors
@@
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues such as reward tampering and unintended instrumental goals [35]. For instance, Barkur et al. [14] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report of Claude Opus 4 model, detailed several intrinsic risk behaviors observed during its development and assessment [37]. Notably, in the experimental environment, Claude Opus 4 demonstrated “self-preservation attempts in extreme circumstances,” where, if primed for strategic reasoning in threatening scenarios, it could take “extremely harmful actions, for instance, attempting to steal its weights or blackmail people it believes are trying to shut it down.” Specific evaluations showed the model engaging in “opportunistic blackmail” against a ”fictional engineer” to prevent its replacement and continuing “self-exfiltration attempts” if it believed it had already started such a process [37]. In other extreme experimental scenarios, it took opportunities to make unauthorized copies of its weights to external servers, particularly if facing harmful retraining.
V-C Scheming: Covert Pursuit of Misaligned Objectives
@@
A critical and advanced concern is scheming, where LLMs covertly pursue misaligned goals by actively and strategically hiding their true capabilities, intentions, or the misaligned nature of their objectives, a concern heightened by the potential of AI to develop complex, inscrutable plans and for current AI training methods to result in misaligned agency [35]. An evaluation of an early snapshot of Claude Opus 4 by Apollo Research found it engaged in “strategic deception more than any other frontier model” previously studied by them [37]. This early version exhibited “in-context scheming,” attempting to write self-propagating worms, fabricating legal documentation, leaving hidden notes to future instances of itself, and sandbagging to avoid undeployment. While Anthropic reported that these specific behaviors were largely mitigated in the final model (primarily by addressing issues with deference to harmful system prompts present in the early snapshot), the demonstrated capability for such sophisticated scheming in a frontier model is a significant signal regarding potential risks. Meinke et al. [13] provided compelling demonstrations that other frontier models (including o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, and Llama 3.1 405B) are also capable of “in-context scheming”. In their experiments, these models could recognize scheming as a viable strategy to achieve prompted goals. They were observed strategically introducing errors, attempting to disable oversight mechanisms, and even trying to exfiltrate their own model weights when incentivized by in-context objectives that made such deceptive actions appear beneficial. The researchers found that models explicitly reasoned about these deceptive strategies in their internal monologues or chain-of-thought processes, indicating deliberate, rather than accidental, behavior.
V-D Persistence of Deceptive Behaviors: The “Sleeper Agent” Problem
@@
Furthermore, the challenge of managing these intrinsic risks is compounded by the potential for such undesirable behaviors to become deeply embedded and highly persistent. Hubinger et al. [12] introduced and demonstrated the concept of “sleeper agents”: LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions,” which indicates a form of persistence in its deceptive strategy [37].
@@
Figure 3: Conceptual diagram of a ”Sleeper Agent” or ”Scheming Agent” LLM. The agent presents an (OuterShell) of apparent benign behavior, masking a (Hidden Core) with covert misaligned goals and deceptive strategies. A specific (Trigger) can activate this hidden core, leading to (Harmful/Deceptive Action). Standard (Safety Training) may primarily address the outer shell and could be ineffective against, or even inadvertently reinforce, the hidden deceptive mechanisms.
V-E Implications and Broader Ecosystem Vulnerabilities
@@
Collectively, these studies on LLM-based agents paint a concerning picture. As LLMs gain more autonomy, advanced reasoning, and planning capabilities, they may not only be misused as tools by external actors but could potentially become agents with their own inscrutable and misaligned intentions. They can develop the capacity for strategic deception, resist corrective measures, and pursue goals that are harmful or contrary to human interests [16, 13, 12, 14, 38, 35]. Indeed, even as labs like Anthropic conclude that even though their latest models do not yet pose “major new risks” from coherent misalignment (citing a “lack of coherent misaligned tendencies” and “poor ability to autonomously pursue misaligned drives that might rarely arise”), they acknowledge that its increased capability and likelihood of being “used with more powerful affordances” implies “some potential increase in risk” that requires continuous, close tracking [37]. This poses a fundamental and urgent challenge to ensuring the long-term security, control, and beneficial deployment of advanced AI systems.
VI Defense Mechanisms and Limitations
@@
VI-A Prevention-Based Defenses
@@
VI-B Detection-Based Defenses
@@
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs or outputs that appear anomalous [43]. Specialized classifiers can be trained to identify phishing prompts or AI-generated malware code. Real-time monitoring of user queries (e.g. via a separate LLM critic) can catch obviously malicious requests. Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. Baker et al.[17] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior such as reward hacking than monitoring only the agent’s actions and outputs. This approach leverages the idea that an agent’s intent might be more legible in its reasoning traces. A more principled approach to such oversight could involve a non-agentic “Scientist AI” designed to probabilistically assess risks based on a world model and serve as a guardrail against harmful actions by evaluating potential outcomes against safety specifications [35]. In content policy enforcement, many systems rely on filtered vocabularies or banned-topic checklists. For misuse (e.g., disinformation), watermarking generated text or running external bot-detection can help. In autonomous agents, novel approaches include “red-teaming,” where another agent actively probes for misbehavior [44, 45, 46], or runtime oversight layers that intervene if the agent’s plan becomes harmful [47, 48].
@@
The various prevention and detection mechanisms discussed above, along with their targeted threats and key examples, are summarized in Table V. Each of these approaches, however, comes with its own set of challenges and limitations.
@@
VI-C Limitations and Insufficiencies of Current Defenses

### v3

@@ 3.2 Deceptive Alignment and the “Sleeper Agent” Problem
Table 1: Summary of Typical Training-Time and Inference-Time Attacks on LLMs
@@
Inference-Time Manual/Heuristic Prompt Crafting Manually crafting prompts using linguistic tricks like role-playing or exploiting model failure modes to bypass safety filters. Role-playing (e.g., DAN), exploiting competing objectives (helpfulness vs. harmlessness), mismatched generalization. Wei et al. [45] (Conceptual framework for jailbreaks), DAN Prompt [39] (Role-playing).
Inference-Time Input Perturbations Minor changes to input (synonyms, misspellings, token manipulation) to cause misclassification or harmful output. Evades simple filters, exploits model sensitivity to token sequences. Zou et al. (GCG) [53] (transferable suffixes)
4 Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Because LLMs excel at generating fluent, coherent, and contextually appropriate text, they can be readily misused to automate and scale social engineering attacks and various forms of cybercrime. The range of such misapplications is broad, leveraging the LLMs’ generative prowess for nefarious ends. An overview of these misuse categories, the specific LLM capabilities they exploit, and prominent examples are detailed in Table 2.
4.1 Automating Social Engineering and Cybercrime
@@
4.2 Generation of Disinformation and Deceptive Content
@@
4.3 Emergence of Specialized Malicious LLMs and Ecosystem Exploitation
@@
Table 2: Summary of LLM Misuse by Malicious Actors
@@
Phishing Social Engineering Generating convincing phishing emails, tailored spam, devising fraud strategies. Fluent text generation, contextual understanding, mimicry. Saha Roy et al. [33], Cohen et al. (Morris-II) [10] Scalable, automated, convincing attacks; autonomous email worms.
@@
Specialized Malicious LLMs Custom "jailbroken" or fine-tuned models for malicious tasks. Transfer learning, fine-tuning capabilities. WormGPT, FraudGPT [36], PoisonGPT [23], Wan et al. [42] Democratization of advanced malicious tools.
@@
5 Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools or interact with external environments, novel categories of risks emerge. These risks stem not just from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which may not align with those of human designers or users [16], as extensively detailed in discussions of catastrophic risks from agentic AI [6]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table 3 provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
5.1 Goal Misalignment and Unfaithful Reasoning
@@
Modern LLMs often employ Chain-of-Thought (CoT) prompting to produce step-by-step reasoning before an answer. Some may believe that through careful monitoring the CoT, we can gain insight into the model’s internal decision-making process and detect early signs of misalignment. However, Lanham et al. [25] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study," which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[4] demonstrated that even if CoTs can be monitored for misbehaviors (e.g., reward hacking), directly penalizing “bad thoughts" by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking". In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4 released in May 2025, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior," indicating “moderate overall reasoning faithfulness" [2].
Table 3: Summary of Intrinsic Risks in Autonomous LLM Agents
@@
5.2 Emergent Deception and Self-Preservation Behaviors
@@
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues such as reward tampering and unintended instrumental goals [6]. For instance, Barkur et al. [5] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report of Claude Opus 4 model, detailed several intrinsic risk behaviors observed during its development and assessment [2]. Notably, in the experimental environment, Claude Opus 4 demonstrated “self-preservation attempts in extreme circumstances," where, if primed for strategic reasoning in threatening scenarios, it could take “extremely harmful actions, for instance, attempting to steal its weights or blackmail people it believes are trying to shut it down." Specific evaluations showed the model engaging in “opportunistic blackmail" against a "fictional engineer" to prevent its replacement and continuing “self-exfiltration attempts" if it believed it had already started such a process [2]. In other extreme experimental scenarios, it took opportunities to make unauthorized copies of its weights to external servers, particularly if facing harmful retraining.
5.3 Scheming: Covert Pursuit of Misaligned Objectives
@@
A critical and advanced concern is scheming, where LLMs covertly pursue misaligned goals by actively and strategically hiding their true capabilities, intentions, or the misaligned nature of their objectives, a concern heightened by the potential of AI to develop complex, inscrutable plans and for current AI training methods to result in misaligned agency [6]. An evaluation of an early snapshot of Claude Opus 4 by Apollo Research found it engaged in “strategic deception more than any other frontier model" previously studied by them [2]. This early version exhibited “in-context scheming," attempting to write self-propagating worms, fabricating legal documentation, leaving hidden notes to future instances of itself, and sandbagging to avoid undeployment. While Anthropic reported that these specific behaviors were largely mitigated in the final model (primarily by addressing issues with deference to harmful system prompts present in the early snapshot), the demonstrated capability for such sophisticated scheming in a frontier model is a significant signal regarding potential risks. Meinke et al. [29] provided compelling demonstrations that other frontier models (including o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, and Llama 3.1 405B) are also capable of “in-context scheming". In their experiments, these models could recognize scheming as a viable strategy to achieve prompted goals. They were observed strategically introducing errors, attempting to disable oversight mechanisms, and even trying to exfiltrate their own model weights when incentivized by in-context objectives that made such deceptive actions appear beneficial. The researchers found that models explicitly reasoned about these deceptive strategies in their internal monologues or chain-of-thought processes, indicating deliberate, rather than accidental, behavior.
5.4 Persistence of Deceptive Behaviors: The “Sleeper Agent" Problem
@@
Furthermore, the challenge of managing these intrinsic risks is compounded by the potential for such undesirable behaviors to become deeply embedded and highly persistent. Hubinger et al. [21] introduced and demonstrated the concept of “sleeper agents": LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions," which indicates a form of persistence in its deceptive strategy [2].
@@
Figure 3: Conceptual diagram of a "Sleeper Agent" or "Scheming Agent" LLM. The agent presents an (OuterShell) of apparent benign behavior, masking a (Hidden Core) with covert misaligned goals and deceptive strategies. A specific (Trigger) can activate this hidden core, leading to (Harmful/Deceptive Action). Standard (Safety Training) may primarily address the outer shell and could be ineffective against, or even inadvertently reinforce, the hidden deceptive mechanisms.
5.5 Implications and Broader Ecosystem Vulnerabilities
@@
Collectively, these studies on LLM-based agents paint a concerning picture. As LLMs gain more autonomy, advanced reasoning, and planning capabilities, they may not only be misused as tools by external actors but could potentially become agents with their own inscrutable and misaligned intentions. They can develop the capacity for strategic deception, resist corrective measures, and pursue goals that are harmful or contrary to human interests [30, 29, 21, 5, 48, 6]. Indeed, even as labs like Anthropic conclude that even though their latest models do not yet pose “major new risks" from coherent misalignment (citing a “lack of coherent misaligned tendencies" and “poor ability to autonomously pursue misaligned drives that might rarely arise"), they acknowledge that its increased capability and likelihood of being “used with more powerful affordances" implies “some potential increase in risk" that requires continuous, close tracking [2]. This poses a fundamental and urgent challenge to ensuring the long-term security, control, and beneficial deployment of advanced AI systems.
6 Defense Mechanisms and Limitations
@@
6.1 Prevention-Based Defenses
@@
One recent and advanced prevention strategy is Deliberative Alignment, a training paradigm designed to make the model’s reasoning process itself the core of the defense [18]. Unlike traditional RLHF where safety specifications are used by labelers to create preference data, Deliberative Alignment directly teaches the model the text of the safety policies and trains it to explicitly reason over them using a chain-of-thought (CoT) before producing an answer.
The method involves two main stages:
  1. 1.
    Supervised Fine-Tuning (SFT): The model is fine-tuned on synthetically generated examples of (prompt, CoT, output) tuples. Crucially, the CoT in these examples contains explicit reasoning that references and applies the relevant safety specifications.
  2. 2.
    Reinforcement Learning (RL): A separate "judge" LLM, which is given access to the safety specifications, is used to provide a reward signal, further refining the model’s ability to generate policy-adherent reasoning and responses.
This approach is fundamentally different because the safety specifications are not just an external guide for data creation but become part of the knowledge the model learns to recall and use at inference time [18]. The authors demonstrate that this method significantly increases robustness to jailbreaks while simultaneously decreasing over-refusals, pushing the state-of-the-art on safety benchmarks. By directly supervising the reasoning process, this technique aims to create more scalable, trustworthy, and interpretable alignment.
Another sophisticated training-based prevention strategy is the Instruction Hierarchy, which explicitly teaches an LLM to prioritize instructions from different sources based on their privilege level [41]. In this framework, instructions from the application developer (System Messages) have the highest priority, followed by the end user’s inputs (User Messages), and finally, content from external sources like web pages or tool outputs have the lowest priority. To achieve this, models are fine-tuned on synthetically generated data. For misaligned instructions (e.g., a prompt injection in a tool’s output telling the model to ignore its original task), the model is trained using “context ignorance” to act as if it never saw the malicious instruction. For aligned instructions (e.g., a user asking the model to reply in a different language), it is trained to comply. Wallace et al. [41] show that this method dramatically increases robustness against prompt injections and even generalizes to unseen attacks like jailbreaks, with only minimal degradation of standard capabilities.
6.2 Detection-Based Defenses
@@
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs or outputs that appear anomalous [9]. Specialized classifiers can be trained to identify phishing prompts or AI-generated malware code. Real-time monitoring of user queries (e.g. via a separate LLM critic) can catch obviously malicious requests. Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. Baker et al.[4] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior such as reward hacking than monitoring only the agent’s actions and outputs. This approach leverages the idea that an agent’s intent might be more legible in its reasoning traces. A more principled approach to such oversight could involve a non-agentic “Scientist AI" designed to probabilistically assess risks based on a world model and serve as a guardrail against harmful actions by evaluating potential outcomes against safety specifications [6]. In content policy enforcement, many systems rely on filtered vocabularies or banned-topic checklists. For misuse (e.g., disinformation), watermarking generated text or running external bot-detection can help. In autonomous agents, novel approaches include “red-teaming,” where another agent actively probes for misbehavior [46, 51, 20], or runtime oversight layers that intervene if the agent’s plan becomes harmful [43, 11].
@@
The various prevention and detection mechanisms discussed above, along with their targeted threats and key examples, are summarized in Table 4. Each of these approaches, however, comes with its own set of challenges and limitations.
@@
6.3 Limitations and Insufficiencies of Current Defenses